Demand in a market dominated by two firms (a Cournot duopoly) is determined according to: P = 300 – 4(Q1 + Q2), where P is the market price, Q1 is the quantity demanded by Firm 1, and Q2 is the quantity demanded by Firm 2. The marginal cost and average cost for each firm is constant; AC=MC = $76. The cournot-duopoly equilibrium quantity produced by each firm is _____. Hint: Write your answer to two decimal plac
Answer #1

Competitive output: P = MC

P = 300 - 4Q

MC =76

P = MC

300 - 4Q = 76

Q = 224/4

56

Cournot output = (2/3)*Competitive output

= (2/3)*56

= 37.33

Each firm produces= 18.66